# STaR: Seamless Spatial-Temporal Aware Motion Retargeting with Penetration and Consistency Constraints

### v1

@@ Abstract
Motion retargeting seeks to faithfully replicate the spatio-temporal motion characteristics of a source character onto a target character with a different body shape. Apart from motion semantics preservation, ensuring geometric plausibility and maintaining temporal consistency are also crucial for effective motion retargeting. However, many existing methods prioritize either geometric plausibility or temporal consistency. Neglecting geometric plausibility results in interpenetration while neglecting temporal consistency leads to motion jitter. In this paper, we propose a novel sequence-to-sequence model for seamless Spatial-Temporal aware motion Retargeting (STaR), with penetration and consistency constraints. STaR consists of two modules: (1) a spatial module that incorporates dense shape representation and a novel limb penetration constraint to ensure geometric plausibility while preserving motion semantics, and (2) a temporal module that utilizes a temporal transformer and a novel temporal consistency constraint to predict the entire motion sequence at once while enforcing multi-level trajectory smoothness. The seamless combination of the two modules helps us achieve a good balance between the semantic, geometric, and temporal targets. Extensive experiments on the Mixamo and ScanRet datasets demonstrate that our method produces plausible and coherent motions while significantly reducing interpenetration rates compared with other approaches.
@@ 1 Introduction
Motion retargeting is a fundamental task in computer vision and graphics. It involves transferring motion from one character to another—preserving the source character’s dynamics and style while adapting to the target character’s unique structure and geometric constraints. It is widely employed in film production, gaming, and robotic animation. With the rise of the digital avatars in virtual environments, rapid advancements in motion generation and digital avatars have further intensified research interest in motion retargeting, which bridges motion capture and character animation. As the demand for authentic visual effects grows, expectations for generated motion have become more stringent. Issues like geometric penetration and unnatural temporal jitter can significantly undermine realism.
@@ 3.1 Spatio-Temporal Aware Motion Retargeting
As shown in Fig. 2, given a source motion $\mathbf{m}_{A}$ from the source character $A$ and the T-posed meshes of both the source character ${mesh}_{A}$, and target character ${mesh}_{B}$, the goal is to generate the target motion $\mathbf{m}_{B}$ while preserving motion semantics, geometric plausibility, and temporal consistency. To achieve this, we decompose the input $\mathbf{m}_{A}$ into skeletal motion, represented by rotation quaternions $\mathbf{Q}\in\mathbb{R}^{T\times K\times 4}$, and global movement $\mathbf{D}\in\mathbb{R}^{T\times 4}$, capturing both local skeletal rotations and the root joint’s trajectory. Here, $T$ denotes the length of the motion sequence, and $K$ represents the number of skeletal joints. To account for height differences between the source and target characters, the global motion is normalized and later de-normalized.
@@
Figure 3: Details of STaR. (a) The detailed architecture of the spatio-temporal model. (b) The constraints on skeletal motion, including temporal consistency loss and semantic loss. (c) The limb penetration loss, which is applied on the skinned motion.
Motion Encoder and Shape Encoder. As shown in Fig. 3 (a), like most retargeting methods [11, 20, 19], we utilize a motion encoder to process the skeletal motion $\mathbf{Q}_{A}$ to motion feature $\mathbf{f}^{q}\in\mathbb{R}^{T\times K\times C}$, where $C$ is the latent length. However, this sparse input cannot provide dense information for motion retargeting, when there is dense supervision signals applied.
To address this, we incorporate a frozen off-the-shelf point cloud network as a geometric feature extractor. Since most point cloud learning methods require a fixed number of input points, we randomly sample point sets from the T-posed meshes of both source and target characters at each iteration. Because the sampled points vary in each iteration, the network is exposed to different geometric subsets over time, forcing it to develop a more robust shape awareness rather than overfitting to specific sampled regions. This strategy enhances the network’s ability to generalize across different body shapes while ensuring a consistent understanding of detailed geometry.
@@ 3.2 Limb Penetration Constraint Module
To prevent penetration, we propose limb penetration loss. As illustrated in Fig. 4 (a), this loss is designed based on the observation that penetration occurs not only between the limbs and the main body [20], but also between the limbs. To comprehensively mitigate penetration, we penalize each limb’s penetration with all other body parts, as illustrated in Fig. 3 (c).
The calculation of the limb penetration loss is inspired by Chamfer Distance and point cloud registration [4], which are based on measuring the distance between two point sets. As clarified in Fig. 4, the loss calculation follows these steps (1) deforming the T-posed vertices and normal vectors of the target character according to $Q_{B}$ using Linear Blend Skinning (LBS) to obtain posed vertices for the limb and the rest of the body, serving as query and reference vertices, respectively; (2) computing the Chamfer Distance between the query and reference vertices; identifying the nearest reference vertex for each query vertex; and (4) multiplying the vector from the query vertex to its nearest reference vertex by the deformed normal vector of the reference vertex, yielding the penetration loss for that query vertex. Specifically, for each query vertex $\mathbf{e}$ on the limbs, its nearest reference vertex $\mathbf{e}^{r}$ with normal $\mathbf{n}^{r}$ is found by function $\eta(\cdot)$. The SDF value for the query vertex $\phi(\mathbf{e})$ is calculated by the dot product $\mathbf{v}^{r}\cdot\mathbf{n}^{r}$, where $\mathbf{v}^{r}=\mathbf{e}-\mathbf{e}^{r}$. The overall penetration penalty $\mathcal{L}_{sdf}$ is defined as follows:
@@
Please refer to Sec. A in the supplemental materials for the detailed image which displays the conditions of both penetrating vertices and non-penetrating ones.
Figure 4: Illustration of different types of interpenetration. Existing methods [20] focus on the interpenetration between limbs and body or head. While for motion like clapping, two hands tend to penetrate each other. We design a novel normal-based SDF loss to include all kinds of interpenetration.
@@ 3.3 Temporal Consistency Constraint Module
Simple smoothness losses used by other works focus on consecutive frames, whose receptive field is quite limited and fails to perceive the smoothness of the full motion path. To narrow the consistency gap between the generated motion and the source motion, we introduce multi-level temporal consistency loss. We first obtain the motion trajectory of each joint by manipulating the skeleton with Forward Kinematics (FK) according to the predicted skeletal rotations $\mathbf{Q}_{B}$, as demonstrated in Fig. 3 (b). Unlike [19], we do not normalize the motion by the character height or skeleton length. Instead, we normalize the trajectory by itself, scaling it to a unit cube to neglect the influence of different skeleton structure, as displayed in Fig. 5 (b). We maintain the motion dynamics by constraining the pair-wise motion vectors. For the joint $k$, the normalized motion trajectory is $\{\mathbf{c}^{k}_{t}\}_{t=1}^{T}$. For each point pair $(\mathbf{c}^{k}_{i},\mathbf{c}^{k}_{j}),i,j\in\{1,...T\}$, we calculate the motion vector $\mathbf{m}_{i,j}^{k}$ from the start point $\mathbf{c}^{k}_{i}$ to the endpoint $\mathbf{c}^{k}_{j}$. Stacking the motion vectors together, we create a motion matrix $\mathbf{M}^{k}\!\in\mathbb{R}^{T\times T\times 3}$. As the Fig. 5 (b) shows, we use the difference of the motion matrix $\mathbf{M}^{k}$ in the temporal consistency loss. The motion matrix is calculated for each joint of the source motion, and it acts as the supervision of the motion dynamics of its correspondent joint in the retargeted motion. The temporal consistency loss is defined as:
$\mathcal{L}_{t}=\frac{1}{K}\sum_{k=1}^{K}\left\lVert\mathbf{M}_{A}^{k}-\mathbf{M}_{B}^{k}\right\rVert^{2}_{2}.$ (3)
@@ 3.4 Loss Functions
As illustrated in Fig. 3, unlike previous multi-stage models [20, 19], we train STaR with an end-to-end manner. And as there are no paired datasets for the motion retargeting task, we exploit as many supervision signals as possible from the source motion to achieve unsupervised learning. Apart from the limb penetration loss and temporal consistency loss introduced in Sec. 3.2 and Sec. 3.3, we utilize several semantic losses to preserve the motion semantics.
Reconstruction Loss. We adopt a self-reconstruction strategy following [11, 20]. The objective of self-reconstruction is to constrain the retargeting of the source motion $\mathbf{Q}_{A}$ onto the source character $A$. This strategy assists the network in searching the static points of the neural function, and ensuring the stabilization of the model during training. The reconstruction loss $\mathcal{L}_{rec}$ includes skeletal rotation reconstruction and joint position reconstruction, which is defined as:
@@
Constrain Loss. In contrast to the reconstruction loss, while retargeting motion to the target character $B$, we use the constrain loss to limit modification while reducing penetration. To constrain the modification as small as possible, we encourage the retargeted motion $\mathbf{Q}_{B}$ to be close to the source motion $\mathbf{Q}_{A}$:
@@ 4 Experiments
Datasets. We train and evaluate STaR, along with comparative methods, on the Mixamo dataset [2], a collection of diverse motion clips performed by artificial characters with various skeleton configurations and body shapes. We collect 2,400 motion instances from 11 characters, splitting them into training and testing sets. The training set consists of 1,905 motion sequences, with source and target characters sampled from a pool of 7 characters per sequence. Each source motion is randomly sampled for 60 frames to facilitate motion retargeting training. For testing, we follow the evaluation setup of previous works [16, 11, 1, 20]. The test set includes both Seen and Unseen Characters (SC, UC), as well as Seen and Unseen Motion clips (SM, UM), forming four subsets: SC+SM, UC+SM, SC+UM, and UC+UM, where only SC+SM is encountered during training. We sample 100 pairs per subset, segment each motion into 120-frame sequences, and report final results as the average across all subsets. To assess performance on real-human motion retargeting, we also test on the ScanRet dataset [18], which contains 8,298 motion clips from 100 human actors. During inference, we follow the official split from [18], consisting of 9 motion clips and 10 human actors.
Implementation details. We select $K=22$ nodes per character and sample $N=1024$ vertices for geometric embedding in each forward pass of the network. The feature length is set to $S=64$. The hyperparameters are configured as follows: $\lambda_{rec}=0.1$, $\lambda_{con}=0.1$, $\lambda_{lp}=5.0$, $\lambda_{tc}=1.0$, and $\lambda_{j}=1.0$. The limb penetration loss is partially implemented with CUDA [13] for acceleration. We set $N_{l}$ and $N_{l}^{r}$ to $(50,100,200,400)$, and $(500,1,000,2,000,4,000)$ respectively, and the corresponding results can be found in Sec. 4.3. We train the network using the Adam optimizer [8] and incorporate a frozen point cloud model from PCT [7]. The entire training process runs for 50 epochs.
@@ 4.1 Qualitative Results
We compare our method with “Copy,” SAN [1], R2ET [20], and MeshRet [18], and visualize the skinned results in Fig. 6. The first row shows that when the source and target characters have similar body shapes, all methods preserve motion semantics well, but STaR produces results closest to the source motion. In the second and third rows, when retargeting motion to characters with significantly different body shapes, penetration issues become more pronounced. Copy and SAN [1] struggle to adapt motion naturally when transferring motion from slim to fat, leading to severe penetration. MeshRet [18], though designed for geometric constraints, fails when the source motion already contains interpenetration. R2ET produces competitive results but suffers from severe motion jitter. Despite body shape constraints, STaR effectively preserves motion semantics while significantly reducing penetration artifacts through full-body geometric correction. The last row, which is the preceding motion frame relative to the third row, partially reflects temporal consistency, showing that STaR produces a more stable and natural motion trajectory. Unlike other methods that aligned frame-wise arm movements, STaR maintains a larger motion space and refines motion trajectories even when no interpenetration occurs, demonstrating its ability to adjust trajectories across the entire sequence. We further infer STaR on real-human dataset, ScanRet [18], and the results are shown in Sec. C in supplemental materials.
@@ 4.2 Quantitative Results
Table 1 presents the comparison between our method and state-of-the-art approaches. The “Copy” baseline achieves the lowest MSE and local MSE, as Mixamo [2] may copy and modify motion to create the dataset. While this works well for characters with similar skeletal structures, it often disrupts motion semantics and increases interpenetration for characters with different body shapes Fig. 6. Among skeletal motion retargeting methods, R2ET [20] achieves the lowest MSE and local MSE, demonstrating its ability to preserve motion semantics. Our method, also correction-based, maintains a low MSE while adjusting metric weights to prioritize reduced penetration rate and curvature.
Regarding geometric plausibility, skeletal motion retargeting methods (PM-Net [11], SAN [1]) exhibit penetration rates of around $11.30\%$, as they do not account for body geometry. Methods incorporating geometric constraints, such as R2ET [20] and MeshRet [18], reduce penetration but introduce excessive motion jitter, reflected in high curvature. Our method, STaR, achieves the lowest penetration rate ($7.99\%$), outperforming R2ET [20] by $16.4\%$ while also reducing curvature by $70.1\%$ (10.66 vs. 35.67), with only a slight increase in MSE. This demonstrates that STaR effectively balances motion semantics, geometric plausibility, and temporal consistency, highlighting the benefits of our spatio-temporal model, limb penetration constraint, and temporal consistency constraint.
We also apply a temporal filter to R2ET [20] to mitigate temporal jitter, resulting in a significant reduction in curvature (from 35.67 to 11.15). However, this post-processing step does not improve other metrics, such as penetration rate or MSE, as it does not fundamentally enhance the model’s ability to generate spatially plausible motion.
For human perception evaluation, please refer to our user study results in Sec. D.
Table 1: Quantitative comparisons with state-of-the-art methods. MSElc is the local MSE. R2ET [20] w/ Filter refers to applying a temporal filter to the results of R2ET [20]. Lower values indicate better performance for all metrics.
@@
STaR 0.0368 0.0174 7.99 10.66

### v2

@@ Abstract
Motion retargeting seeks to faithfully replicate the spatio-temporal motion characteristics of a source character onto a target character with a different body shape. Apart from motion semantics preservation, ensuring geometric plausibility and maintaining temporal consistency are also crucial for effective motion retargeting. However, many existing methods prioritize either geometric plausibility or temporal consistency. Neglecting geometric plausibility results in interpenetration, while neglecting temporal consistency leads to motion jitter. In this paper, we propose a novel sequence-to-sequence model for seamless Spatial-Temporal aware motion Retargeting (STaR), with penetration and consistency constraints. STaR consists of two modules: (1) a spatial module that incorporates dense shape representation and a novel limb penetration constraint to ensure geometric plausibility while preserving motion semantics, and (2) a temporal module that utilizes a temporal transformer and a novel temporal consistency constraint to predict the entire motion sequence at once while enforcing multi-level trajectory smoothness. The seamless combination of the two modules helps us achieve a good balance between the semantic, geometric, and temporal targets. Extensive experiments on the Mixamo and ScanRet datasets demonstrate that our method produces plausible and coherent motions while significantly reducing interpenetration rates compared with other approaches. Code page: https://github.com/XiaohangYang829/STaR.
footnotetext: For the purpose of open access, the author has applied a Creative Commons Attribution (CC-BY) license to any Author Accepted Manuscript version arising.
@@ 1 Introduction
Motion retargeting is a fundamental task in computer vision and graphics. It involves transferring motion from one character to another—preserving the source character’s dynamics and style while adapting to the target character’s unique structure and geometric constraints. It is widely employed in film production, gaming, and robotic animation. With the rise of digital avatars in virtual environments, rapid advancements in motion generation and digital avatars have further intensified research interest in motion retargeting, which bridges motion capture and character animation. As the demand for authentic visual effects grows, expectations for generated motion have become more stringent. Issues like geometric penetration and unnatural temporal jitter can significantly undermine realism.
@@ 3.1 Spatio-Temporal Aware Motion Retargeting
As shown in Fig. 2, given a source motion $\mathbf{m}_{A}$ from the source character $A$ and the T-posed meshes of the source character ${mesh}_{A}$, and target character ${mesh}_{B}$, the goal is to generate the target motion $\mathbf{m}_{B}$ while preserving motion semantics, geometric plausibility, and temporal consistency. To achieve this, we decompose the input $\mathbf{m}_{A}$ into skeletal motion, represented by rotation quaternions $\mathbf{Q}\in\mathbb{R}^{T\times K\times 4}$, and global movement $\mathbf{D}\in\mathbb{R}^{T\times 4}$, capturing both local skeletal rotations and the root joint’s trajectory. Here, $T$ denotes the length of the motion sequence, and $K$ represents the number of skeletal joints. Because characters vary in height, we compensate for this by normalizing global motion, then restoring its original scale afterward.
@@
Figure 3: Details of STaR. (a) The detailed architecture of the spatio-temporal model. (b) The constraints on skeletal motion include temporal consistency loss and semantic loss. (c) The limb penetration loss is applied to the skinned motion.
Motion Encoder and Shape Encoder. As shown in Fig. 3 (a), like most retargeting methods [11, 20, 19], we utilize a motion encoder to process the skeletal motion $\mathbf{Q}_{A}$ to motion feature $\mathbf{f}^{q}\in\mathbb{R}^{T\times K\times C}$, where $C$ is the latent length. However, this sparse input cannot provide dense information for motion retargeting when there are dense supervision signals applied.
To address this, we incorporate a frozen off-the-shelf point cloud network [7] as a geometric feature extractor. Since most point cloud learning methods require a fixed number of input points, we randomly sample point sets from the T-posed meshes of both source and target characters at each iteration. Because the sampled points vary in each iteration, the network is exposed to different geometric subsets over time, forcing it to develop a more robust shape awareness rather than overfitting to specific sampled regions. This strategy enhances the network’s ability to generalize across different body shapes while ensuring a consistent understanding of detailed geometry.
@@ 3.2 Limb Penetration Constraint Module
To prevent penetration, we propose limb penetration loss. As illustrated in Fig. 4 (a), this loss is designed based on the observation that penetration occurs not only between the limbs and the main body [20], but also between the limbs. To mitigate penetration, we penalize each limb’s penetration with all other body parts, as illustrated in Fig. 3 (c).
The calculation of the limb penetration loss is inspired by Chamfer Distance and point cloud registration [4], which are based on measuring the distance between two point sets. As clarified in Fig. 4, the loss calculation follows these steps (1) deforming the T-posed vertices and normal vectors of the target character according to $Q_{B}$ using Linear Blend Skinning (LBS) to obtain posed vertices for the limb and the rest of the body, serving as query and reference vertices, respectively; (2) computing the Chamfer Distance between the query and reference vertices; (3) identifying the nearest reference vertex for each query vertex; and (4) multiplying the vector from the query vertex to its nearest reference vertex by the deformed normal vector of the reference vertex, yielding the penetration loss for that query vertex. Specifically, for each query vertex $\mathbf{e}$ located on the limbs, the nearest reference vertex $\mathbf{e}^{r}$ with normal vector $\mathbf{n}^{r}$ is identified using a Nearest Neighbor Search (NNS) algorithm. The SDF value for the query vertex $\phi(\mathbf{e})$ is calculated by the dot product $\mathbf{v}^{r}\cdot\mathbf{n}^{r}$, where $\mathbf{v}^{r}=\mathbf{e}-\mathbf{e}^{r}$. The overall penetration penalty $\mathcal{L}_{lp}$ is defined as follows:
@@
Please refer to Sec. A in the supplemental materials for the detailed image, which displays the conditions of both penetrating vertices and non-penetrating ones.
Figure 4: Illustration of different types of interpenetration. Existing methods [20] focus on the interpenetration between limbs and body or head. While for motions like clapping, two hands tend to penetrate each other. We design a novel normal-based SDF loss to include all kinds of interpenetration.
@@ 3.3 Temporal Consistency Constraint Module
Simple smoothness losses used by other works focus on consecutive frames, whose receptive field is quite limited and fails to perceive the smoothness of the full motion path. To narrow the consistency gap between the generated motion and the source motion, we introduce a multi-level temporal consistency loss. We obtain the motion trajectory of each joint by manipulating the skeleton with Forward Kinematics (FK) according to the predicted skeletal rotations $\mathbf{Q}_{B}$, as demonstrated in Fig. 3 (b). Unlike [19], we do not normalize the motion by the character height or skeleton length. Instead, we normalize the trajectory by itself, scaling it to a unit cube to neglect the influence of different skeleton structures, as displayed in Fig. 5 (b). We maintain the motion dynamics by constraining the pair-wise motion vectors. For the joint $k$, the normalized trajectory is $\{\mathbf{c}^{k}_{t}\}_{t=1}^{T}$. For each point pair $(\mathbf{c}^{k}_{i},\mathbf{c}^{k}_{j}),i,j\in\{1,...T\}$, we calculate the motion vector $\mathbf{m}_{i,j}^{k}$ from the start point $\mathbf{c}^{k}_{i}$ to the endpoint $\mathbf{c}^{k}_{j}$. Stacking the motion vectors together, we create a matrix $\mathbf{M}^{k}\!\in\mathbb{R}^{T\times T\times 3}$. As the Fig. 5 (b) shows, we use the difference of the motion matrix $\mathbf{M}^{k}$ in the temporal consistency loss. The motion matrix is calculated for each joint of the source motion, and it acts as the supervision of the motion dynamics of its corresponding joint in the retargeted motion. The temporal consistency loss is defined as:
$\mathcal{L}_{tc}=\frac{1}{K}\sum_{k=1}^{K}\left\lVert\mathbf{M}_{A}^{k}-\mathbf{M}_{B}^{k}\right\rVert^{2}_{2}.$ (3)
@@ 3.4 Loss Functions
As illustrated in Fig. 3, unlike previous multi-stage models [20, 19], we train STaR in an end-to-end manner. And as there are no paired datasets for the motion retargeting task, we exploit as many supervision signals as possible from the source motion to achieve unsupervised learning. Apart from the limb penetration loss and temporal consistency loss introduced in Sec. 3.2 and Sec. 3.3, we utilize several semantic losses to preserve the motion semantics.
Reconstruction Loss. We adopt a self-reconstruction strategy following [11, 20]. The objective of self-reconstruction is to constrain the retargeting of the source motion $\mathbf{Q}_{A}$ onto the source character $A$. This strategy assists the network in searching for the static points of the neural function and ensures the stabilization of the model during training. The reconstruction loss $\mathcal{L}_{rec}$ includes skeletal rotation reconstruction and joint position reconstruction, which is defined as:
@@
Constraint Loss. In contrast to the reconstruction loss, while retargeting motion to the target character $B$, we use the constraint loss to limit modification while reducing penetration. To constrain the modification as small as possible, we encourage the retargeted motion $\mathbf{Q}_{B}$ to be close to the source motion $\mathbf{Q}_{A}$:
@@ 4 Experiments
Datasets. We train and evaluate STaR, along with comparative methods, on the Mixamo dataset [2], a collection of diverse motion clips performed by artificial characters with various skeleton configurations and body shapes. We collect 2,400 motion instances from 11 characters, splitting them into training and testing sets. The training set comprises 1,905 motion sequences, with source and target characters sampled from a pool of 7 characters per sequence. Each source motion is randomly sampled for 60 frames to facilitate motion retargeting training. For testing, we follow the evaluation setup of previous works [16, 11, 1, 20]. The test set includes both Seen and Unseen Characters (SC, UC), as well as Seen and Unseen Motion clips (SM, UM), forming four subsets: SC+SM, UC+SM, SC+UM, and UC+UM. Only SC+SM is encountered during training. We sample 100 pairs per subset, segment each motion into 120-frame sequences, and report final results as the average across all subsets. To assess performance on real-human motion retargeting, we also test on the ScanRet dataset [18], which contains 8,298 motion clips from 100 human actors. During inference, we follow the official split from [18], consisting of 9 motion clips and 10 human actors.
Implementation details. We select $K=22$ nodes per character and sample $N=1024$ vertices for geometric embedding in each forward pass of the network. The feature length is set to $S=64$. The hyperparameters are configured as follows: $\lambda_{rec}=0.1$, $\lambda_{con}=0.1$, $\lambda_{lp}=5.0$, $\lambda_{tc}=1.0$, and $\lambda_{j}=1.0$. The limb penetration loss is partially implemented with CUDA [13] for acceleration. We set $N_{l}$ and $N_{l}^{r}$ to (50,100,200,400)(50,100,200,400)( 50 , 100 , 200 , 400 ), and (500,1,000,2,000,4,000)(500,1{,}000,2{,}000,4{,}000)( 500 , 1 , 000 , 2 , 000 , 4 , 000 ) respectively, and the corresponding results can be found in Sec. 4.3. We train the network using the Adam optimizer [8] and incorporate a frozen point cloud model from PCT [7]. The entire training process runs for 50 epochs.
@@ 4.1 Qualitative Results
We compare our method with “Copy”, SAN [1], R2ET [20], and MeshRet [18], and visualize the skinned results in Fig. 6. The first row shows that when the source and target characters have similar body shapes, all methods preserve motion semantics well, but STaR produces results closest to the source motion. In the second and third rows, when retargeting motion to characters with significantly different body shapes, penetration issues become more pronounced. Copy and SAN [1] struggle to adapt motion naturally when transferring motion from slim to fat, leading to severe penetration. MeshRet [18], though designed for geometric constraints, fails when the source motion already contains interpenetration. R2ET produces competitive results but suffers from severe motion jitter. Despite body shape constraints, STaR effectively preserves motion semantics while significantly reducing penetration artifacts through full-body geometric correction. The last row, which is the preceding motion frame relative to the third row, partially reflects temporal consistency, showing that STaR produces a more stable and natural motion trajectory. Unlike other methods that aligned frame-wise arm movements, STaR maintains a larger motion space and refines motion trajectories even when no interpenetration occurs, demonstrating its ability to adjust trajectories across the entire sequence. We further infer STaR on the real-human dataset, ScanRet [18], and the results are shown in Sec. C in the supplemental materials.
@@ 4.2 Quantitative Results
Table 1 compares our method with current state-of-the-art approaches. The “Copy” baseline achieves the lowest MSE and local MSE, since Mixamo [2] may generate dataset samples by copying and slightly modifying motions—yielding artificially low error metrics but at the cost of semantic accuracy and increased interpenetration, especially for characters with differing body shapes, as shown in Fig. 6. Among all motion retargeting methods, R2ET [20] achieves the lowest MSE, demonstrating its ability to preserve motion semantics. Our method, STaR, maintains similarly low MSE while adjusting metric weights to prioritize reduced penetration rate and curvature.
[…]
For human perception evaluation, please refer to Sec. D.
Table 1: Quantitative comparisons with state-of-the-art methods. MSElc is the local MSE. R2ET [20] refers to the debugged version of R2ET [20]. STaR (Curv) assigns a higher weight to motion smoothness. For MSE, MSElc, and Pen%, lower values indicate better performance. For Curv, closer to GT’s Curv is better.
[…]
@@
STaR (Curv) 0.0353 0.0158 8.70 7.76
STaR (Ours Final) 0.0368 0.0174 7.99 10.66